OpenStudy (anonymous):

simplify (4)+(2/xy)/(2)

OpenStudy (zzr0ck3r):

2/xy/(2) is like 2/xy/(2/1) flip it and multiple (2/(xy))(1/2)= 1/xy 4 + 1/(xy) now we need common denominator so multiply 4(top and bottom) by xy 4 = 4(xy)/(xy) (4xy)/(xy) + 1/(xy = (4xy+1)/xy

OpenStudy (anonymous):

Thanks!!

OpenStudy (zzr0ck3r):

np